# Faster shortest-path algorithms using the acyclic-connected tree

### v1

@@ Abstract
This paper gives a fixed-parameter linear algorithm for the single-source shortest path problem (SSSP) on directed graphs. The parameter in question is the nesting width, a measure of the extent to which a graph can be represented as a nested collection of graphs. We present a novel directed graph decomposition called the acyclic-connected tree (A-C tree), which breaks the graph into a recursively nested sequence of strongly connected components in topological order. We prove that the A-C tree is optimal in the sense that its width—the size of the largest nested graph—is equal to the nesting width of the graph. We then provide a linear-time algorithm for constructing the A-C tree of any graph. Finally, we show how the A-C tree allows us to construct a simple variant of Dijkstra’s algorithm which achieves a time complexity of $O(e+n\log w)$, where $n$ ($e$) is the number of nodes (arcs) in the graph and $w$ is the nesting width. The idea is to apply the shortest path algorithm separately to each component in the order dictated by the A-C tree. We obtain an asymptotic improvement over Dijkstra’s algorithm: when $w=n$, our algorithm reduces to Dijkstra’s algorithm, but it is faster when $w\in o(n)$, and linear-time for classes of graphs with bounded width, such as directed acyclic graphs.
@@ 1.1 Motivation
Finding the shortest paths from a node in a weighted graph $G$ is one of the oldest and most fundamental algorithmic research problems, whose modern applications range over everything from robotics [35] to route planning [6]. This problem, now called the Single-Source-Shortest-Path Problem (SSSP), originates with the work of Dijkstra [18], whose eponymous algorithm efficiently computes a shortest path tree to all nodes from a given fixed source node $s$ in $G$. Dijkstra’s algorithm (using Fibonacci heaps) runs in time $O(e+n\log n)$, where $n$ and $e$ are the number of nodes and arcs in the graph, and is known to be worst-case optimal for producing a list of nodes in sorted order of distance from the source node [21], a problem called Distance Ordering (DO). In a recent work [28], a new variant of Dijkstra’s algorithm using a ‘working-set’ heap was shown to be universally optimal for DO, a form of beyond-worst-case analysis which measures worst-case performance for each graph individually, up to reweighting of the arcs.
The SSSP problem is somewhat easier than the DO problem, because the nodes do not have to be sorted—Dijkstra’s algorithm also produces a shortest-path tree as well as the sorted order. However, the lower bounds for DO are derived from the problem of comparison sorting. By contrast, SSSP does not share this lower bound—instead, the complexity is only well-understood in special cases (see Section 1.3).
In fact, we know that classes of graphs exist where SSSP is strictly easier than DO. An important example is directed acyclic graphs (DAGs), where SSSP can be solved in linear time [53]. Inspired by [28], we would like to obtain new beyond-worst-case results for SSSP. That is, we want a shortest path algorithm that has the same time complexity as Dijkstra’s algorithm in the worst case, but outperforms Dijkstra’s when the given graph possesses additional structure. Our focus in this paper is on hierarchical structure in particular, that is, when a graph can be hierarchically decomposed into a nested collection of smaller graphs. This is an important property for applications, where the concept of modularity naturally leads to such a nested structure. We show later that nesting collectively generalises several natural graph ideas, including dominance, connected components and acyclicity (see Section 2). In particular, every DAG can be decomposed into a nested collection of graphs with two nodes each.
Example 1.1.
@@
Figure 1: A graph $G$, and a subset of its nodes (Fig. 1(a)). This set of nodes allows $G$ to be decomposed into a pair of graphs $G_{1}$ and $G_{2}$, with $G_{2}$ nested at a node $X$ of $G_{1}$ (Fig. 1(b)).
Consider the graph $G$ in Figure 1 with source node $s$. Here, the set of nodes $S=\{a,d,e,f,g,h\}$ has an interesting property: all arcs into $S$ go to $h$. This observation allows us to run Dijkstra’s algorithm more efficiently on this graph, by breaking it into a composition of two smaller graphs. We can treat this set as being a smaller single-source graph $G_{2}$ (with source $h$) nested within a node $X$ of an ‘outer’ graph $G_{1}$ (Figure 1(b)). Now we can run Dijkstra’s algorithm on $G_{1}$, and when we reach $X$, run Dijkstra’s algorithm recursively on $G_{2}$, beginning at its source $h$. The resultant time complexity is $O(e_{1}+v_{1}\log v_{1}+e_{2}+v_{2}\log v_{2})\leq O(e_{1}+e_{2}+(v_{1}+v_{2})\log\max(v_{1},v_{2}))$, where $v_{i}$ and $e_{i}$ are the number of nodes and arcs in $G_{i}$, respectively. By comparison, naïvely running Dijkstra’s algorithm on $G$ results in the larger running time of $O(e_{1}+e_{2}+(v_{1}+v_{2})\log(v_{1}+v_{2}))$. The complexity of Dijkstra’s algorithm is determined by the nonlinear heap operations—by splitting the problem into two subgraphs, we get two smaller heaps, therefore, the running time improves.
The key observation in Example 1.1 can be generalised to more than one layer of nesting. Indeed, deferring formal details to later sections, given a graph $G$ decomposed into a nested tree of subgraphs $G_{1},G_{2},\dots,G_{m}$, we can use the same recursive application of Dijkstra’s algorithm to compute the shortest paths in time $O((\sum_{i}e_{i})+(\sum_{i}v_{i})\log\max_{i}v_{i})$.
With this as motivation, we would like to able to algorithmically decompose a graph into a tree of smaller nested subgraphs. Generally, graphs can be decomposed in several distinct ways. We therefore seek a decomposition where the maximum size of the nested subgraphs is minimised, a quantity we call the ‘width’ $w$ of the decomposition. We arrive at the main goal of our paper: obtain beyond-worst-case improvements to SSSP by efficiently constructing a minimal-width decomposition given as a tree of nested graphs.
@@ 1.2 Contribution
In this paper, we present a linear-time algorithm for computing a minimal-width decomposition of a graph $G$ and use this decomposition to solve SSSP in time $O(e+n\log\operatorname{nw}(G))$ using a recursive form of Dijkstra’s algorithm [^1] Our construction allows for any SSSP algorithm to be used here. This means that all improvements on the overall problem (such as [28]), can be immediately extended to graphs of small nesting width.. Here, the parameter $\operatorname{nw}(G)$ is called the nesting width of $G$, defined as the minimal width among all nesting decompositions. Since $\operatorname{nw}(G)\leq n$, this SSSP algorithm is an asymptotic improvement over Dijkstra’s algorithm. In detail, our paper has four main contributions.
First, we formalise a hierarchical decomposition of a single-source graph, which we call a nested decomposition (Definition 2.7). We demonstrate its connection to modules (Definition 2.5), which are subsets of nodes which are reachable only through a single node (cf. Example 1.1). We define the width of a nested decomposition as the maximal size of the subgraphs in the decomposition, and define the nesting width $\operatorname{nw}(G)$ of a graph $G$ as the minimal width over all nested decompositions of $G$ (Definition 2.8).
Second, exploiting the connection with modules, we define a new decomposition on a single-source graph $G$ called the acyclic-connected tree (A-C tree) of $G$ (Definition 3.1). The nodes of the A-C tree are the nodes of $G$, with each node having one incoming arc and possibly several outgoing arcs to disjoint sets of nodes in the tree. See Figure 2(b) for an example. The descendants of a node $a$ are those nodes $b$ where all paths to $b$ contain $a$. The sets of nodes that are children of $a$ define strongly connected components of this subgraph, which are ordered in topological order. The name derives from the fact that this tree defines a set of recursively nested acyclic graphs of strongly connected components. We note that the A-C tree does not depend on the weights of $G$ and prove that it is an optimal decomposition in the sense that it has minimal width.
Third, to use the A-C tree for SSSP, we need to be able to construct it efficiently. Specifically, the construction must be asymptotically faster than $O(e+n\log n)$, which is the complexity of simply running Dijkstra’s algorithm. This is our third contribution: we give a linear-time algorithm, called A-C_tree, for computing the A-C tree. Our approach combines two well-established graph concepts with linear-time algorithms: strongly connected components [53] and dominator trees [4].
Fourth, combining the A-C tree with a recursive application of Dijkstra’s algorithm, we solve SSSP for any graph $G$ in time $O(e+n\log\operatorname{nw}(G))$. This is an asymptotic improvement of Dijkstra’s algorithm. Indeed, in the worst case, $\operatorname{nw}(G)=n\in O(n)$, and our time complexity $O(e+n\log\operatorname{nw}(G))$ coincides with Dijkstra’s algorithm. However, whenever $\operatorname{nw}(G)\in o(n)$, our time complexity is strictly smaller. In particular, if a class of graphs has bounded nesting width (like DAGs), we can solve the shortest path problem in linear time.
@@ 1.3 Related work
Constructing fast shortest path algorithms has been an active area of research since the ‘50s, with pioneering work given by [40, 49, 7, 20, 18]. In particular, [21] showed that Dijkstra’s algorithm [18] (with Fibonacci heaps) is worst-case optimal among all algorithms which visit the nodes in order of distance. If one instead only computes the shortest-path tree, then faster algorithms are possible, depending on assumptions. One approach is to constrain the weights, for example by assuming them to be integers [17, 55, 9, 12] or uniformly bounded [17, 2, 32, 33]. We differ from this line of work since our approach is entirely structural—we assume no restriction on the weights except being non-negative. Other work using structural assumptions include search algorithms for planar graphs [37, 30, 19, 34, 41], minor-free graphs [56], directed acyclic graphs (DAGs) [53], and most recently universal optimality [28]. Generalising the DAG case, [52] showed that SSSP can be solved in time $O(e+n\log k)$ where $k$ is the size of the largest connected component. We obtain this as a special case of our method, because the size of the largest connected component is always at least $\operatorname{nw}(G)-1$, and is often strictly larger. Finally, there are several recent algorithms using continuous optimisation tools [48, 42, 15, 8, 42, 57]. We work instead directly in a discrete setup to be closer to the original (discrete) problem.
@@
This paper introduces the A-C tree as a hierarchical decomposition of a graph. Similar concepts include the low diameter decomposition (see [9, 12] and references within), which groups together strongly connected sets with respect to the weights (essentially, nodes that are close together should be grouped together), with probabilistic guarantees. In particular, [9, 12] also hierarchically extract acyclic parts using topological order between strongly connected sets to compute shortest path trees faster, but only for integral weights (possibly negative integers). Another line of research is the hierarchy-based approach developed by [43, 54, 44, 29]. This approach divides the graph into subgraphs splitting up the original SSSP into smaller SSSPs to search faster. The basic idea is to find subgraphs that are safe to search in greedily (ignoring the remaining graph) up to a certain distance. This idea is quite similar to how the A-C tree divides the SSSP into smaller problems. However, these hierarchy-type approaches ultimately rely on the weights, whereas our decomposition is entirely structural. Further, it is shown in [43] that any hierarchy-type algorithm must run in $\Omega(e+n\log n)$ time in the worst case on a certain family of directed graphs. It is straightforward to show that this family has bounded nesting width, and so our algorithm runs in linear time on these graphs.
@@
The A-C tree proposed in this work uses the notion of a dominator tree, introduced by [45] and with the first algorithm given by [38]. Given a graph with a fixed source node $s$, a node $m$ is said to dominate $v$ if every path (from $s$) to $v$ goes through $m$ (Definition 2.2), and the dominator tree is the tree representing this partial order over all nodes (with root $s$). Efficient algorithms for computing the dominator tree include the classical near-linear time algorithm given by [36], and more sophisticated linear time algorithms [4, 13, 14, 25]. In our paper, the dominator tree essentially defines the ‘acyclic’ part of the A-C tree. In this way, the A-C tree can be seen as an extension of the dominator tree, where we also recursively extract strongly connected components. Dominator trees have been used in several applications, such as the design of efficient compilers [1, 16], modelling ecosystems [3], diagnostic test generation [5], 2-vertex connectivity tests and approximations [23, 24], influencing diffusion networks [26], finding strong bridges in graphs [31], diagnosing memory leaks [39], and computing reachability in graphs subject to constraints [46]. However, dominator trees have rarely been applied to shortest path problems, as we do in this paper. One exception is the recent work of [28] which we discussed in the introduction. There they use the dominator tree in a subroutine to simplify the graph as part of the constructing of a universally optimal algorithm for distance ordering in terms of number of comparison queries. Our usage is similar in flavour, though the dominator plays a much more central role in our definitions and results.
@@
There is a plethora of shortest path methods which rely on a preprocessing step to compute shortest paths faster at run-time [6]. However, the preprocessing step itself is typically computationally expensive. For example, the preprocessing step in [51, 50] is exponential in the number of machines in the hierarchy. Computing the A-C tree in this paper can conceptually be viewed as a preprocessing step, but with only a linear-time overhead.
Finally, the nesting width proposed in this paper is inspired by the use of treewidth [47] in the field of parameterised complexity. Treewidth is defined by ‘tree decompositions’, of which each graph can possibly have many. The width of any such decomposition is a bound for the complexity of many graph algorithms. The treewidth is the minimum width over all tree decompositions. While computing the treewidth is NP-hard, some classes of graphs are known to have bounded treewidth, allowing some otherwise NP-complete problems to be solved on them in polynomial time. In the same way, if a class of graphs has bounded nesting width, we can solve SSSP in linear time. In contrast to treewidth, computing the nesting width can be done very efficiently using the A-C tree.
@@ 1.4 Outline
The remainder of the paper is as follows. Section 2 presents the preliminaries. Section 3 describes the A-C tree. Section 4 proves that the A-C tree is a minimal-width decomposition. Section 5 presents an algorithm for computing the A-C tree and shows that it has linear-time complexity. This algorithm is then used in Section 6 to solve SSSP through recursive application of Dijkstra’s algorithm. An example illustrating the theory is given in Section 7. Finally, Section 8 concludes the paper.

### v2

@@ Abstract
We provide a method to obtain beyond-worst-case time complexity for any single-source-shortest-path (SSSP) algorithm by exploiting modular structures in graphs. The key novelty is a graph decomposition, called the acyclic-connected (A-C) tree, which breaks up a graph into a recursively nested sequence of strongly connected components in topological order. The A-C tree is optimal in the sense that it maximally decomposes the graph, formalised by a parameter called nesting width, measuring the extent to which a graph can be decomposed. We show how to compute the A-C tree in linear time, allowing it to be used as a preprocessing step for SSSP. Indeed, we transform any SSSP algorithm by first computing the A-C tree, and then running the SSSP algorithm in a careful recursive manner on the A-C tree. We illustrate this with two state-of-the-art algorithms: Dijkstra’s algorithm and the recent sparse graph algorithm of Duan et al., obtaining improved time complexities of $O(m+n\log(\operatorname{nw}(G)))$ and $O(m\alpha(n)+m\log^{2/3}(\operatorname{nw}(G)))$, respectively, where $\operatorname{nw}(G)\leq n$ is the nesting width of the graph $G$, and $\alpha(n)$ is the extremely slow-growing inverse Ackermann function. Some classes of graphs, such as directed acyclic graphs, have bounded nesting width, and we obtain linear-time SSSP algorithms for these graphs.
@@ 1.1 Motivation
Finding the shortest paths from a node in a weighted graph is one of the oldest and most fundamental problems in the field of algorithms, whose modern applications range from robotics [36] to route planning [6]. This problem, called the Single-Source-Shortest-Path problem (SSSP), originates from the work of Dijkstra [18], whose eponymous algorithm efficiently computes the tree of shortest paths from a given fixed source node $s$ to all other nodes in a graph $G$. Dijkstra’s original algorithm [18] runs in $O(n^{2})$ steps in an $n$-node graph; the addition of the Fibonacci heap data structure by Fredman and Tarjan [22] improved the complexity of Dijkstra’s algorithm to $O(m+n\log n)$, where $n$ and $m$ are the number of nodes and arcs in the graph, respectively. This matches a known lower bound for Dijkstra’s algorithm: because the nodes are visited in sorted order of distance from the source, the algorithm inherits the $\Omega(n\log n)$ lower bound for comparison sorting.
For many decades, it appeared that the SSSP problem was essentially solved (at least in the comparison-addition model [19]). However, two recent breakthroughs ([29, 19]) have called this into question, and the SSSP is now undergoing a significant revival in theoretical computer science. In particular, as noted in [19], the comparison sorting worst-case lower bound is actually too pessimistic. Instead, SSSP can sometimes—and maybe always—be solved more efficiently than Dijkstra’s algorithm.
Indeed, to solve SSSP, one only needs to return the shortest-path tree. Dijkstra’s algorithm also returns the nodes in sorted order, a strictly harder problem, now called Distance Ordering (DO). The lower bound for DO, derived from the problem of comparison sorting, does not apply to SSSP. This motivates the work of Duan et al. [19], who provide the first algorithm to break the $\Omega(n\log n)$ ‘sorting barrier’ for SSSP. Their algorithm returns the shortest-path tree in $O(m\log^{2/3}n)$ steps, which is strictly smaller than $O(m+n\log n)$ for sparse graphs ($m=\Theta(n)$). This result is the first indication that the worst-case complexity of SSSP may be significantly smaller than that of DO.
In another recent work, Haeupler et al. [29] analysed the beyond-worst-case performance of Dijkstra’s algorithm and found, counterintuitively, that it is not “universally optimal” (even for the DO problem). Universal optimality measures the worst-case performance of an algorithm on a given input graph, for all possible weightings of the arcs. [^1] They then show, equipped with a new data structure called a “working-set heap”, that Dijkstra’s algorithm can be made universally optimal for DO. These two breakthroughs demonstrate that the complexity of SSSP remains open, and, crucially, that new ideas can lead to significant improvement on this fundamental problem.
Our work extends this line of work, beginning with the following observation: despite their strengths, neither of these new algorithms [29, 19] are universally optimal for SSSP. In other words, they do not achieve the worst-case time complexity for SSSP given the graph structure. This is demonstrated by directed acyclic graphs (DAGs), where we know that SSSP can be solved in linear time [55], but both of these state-of-the-art methods take superlinear time (this is clear for [29] since it is lower bounded by $\Omega(n\log n)$, and for [19] see Appendix E).
@@
Figure 1: A graph $G$, and a subset of its nodes (Fig. 1(a)). This set allows $G$ to be decomposed into a pair of graphs $G_{1}$ and $G_{2}$, with $G_{2}$ nested at a node $X$ of $G_{1}$ (Fig. 1(b)).
1.2 Modular structure in graphs
Inspired by [29], we would like to obtain SSSP algorithms that match the worst-case time complexity of state-of-the-art methods (such as [19]), but outperform these methods when the graph possesses additional structure, with the goal of building towards a universally optimaln algorithm for SSSP. [^2] Our method is not universally optimal for SSSP, but it is a necessary step in that direction, because any universally optimal SSSP algorithm must perform at least as well as ours on graphs with bounded nesting width. See Section 8. Our focus in this paper is on modular structure, that is, when a graph can be hierarchically decomposed into a nested collection of smaller graphs. Hierarchical nesting is motivated by the presence of modularity in many real-world applications. It can also be viewed as a generalisation of the DAG case; as we show later, every DAG can be decomposed as a nested collection of two-node graphs. Figure 1 demonstrates such a modular structure in a graph $G$. Here, the set of nodes $S=\{a,d,e,f,g,h\}$ in $G$ has an important property: all arcs into $S$ go to $h$. This allows $G$ to be represted as a pair of graphs $G_{1}$ and $G_{2}$ (Fig. 1(b)), with $G_{2}$ nested in a node $X$ of $G_{1}$. Thus, instead of running a black-box SSSP algorithm on $G$, we can instead run it on $G_{1}$ and then recursively on $G_{2}$ upon reaching $X$, beginning at its source $h$, with the nonlinear component of the time complexity scaling with the size of $G_{1}$ and $G_{2}$. Using Dijkstra’s algorithm as an example, decomposing this graph improves the running time from $O(m_{1}+m_{2}+(n_{1}+n_{2})\log(n_{1}+n_{2}))$ to $O(m_{1}+n_{1}\log n_{1}+m_{2}+n_{2}\log n_{2})\leq O(m_{1}+m_{2}+(n_{1}+n_{2})\log\max(n_{1},n_{2}))$, where $n_{i}$ and $m_{i}$ are the number of nodes and arcs in $G_{i}$, respectively.
@@ 1.3 Contribution
In this paper, we provide a method to obtain beyond-worst-case time complexity for any SSSP algorithm by exploiting modular structures in graphs. The key idea is to first decompose a given graph into a hierarchy of smaller subgraphs and then run the SSSP algorithm on this decomposition. We first formalise the concept of hierarchical decomposition, by defining a nesting decomposition of a graph. In general, a graph typically has several nesting decompositions. We therefore seek one where its width, defined as the maximum size of its nested subgraphs, is minimised. [^3] Intuitively, this corresponds to a graph that is maximally decomposed. This minimal width, called the nesting width (inspired by treewidth [48]), is a novel graph parameter that captures the extent to which a graph can be decomposed. We then introduce a novel combinatorial object called the acyclic-connected (A-C) tree of a graph, which defines how a graph can be broken into a recursively nested sequence of strongly connected components in topological order. The A-C tree is optimal in the sense that it defines a minimal-width nesting decomposition. Crucially, we give a linear-time algorithm for computing the A-C tree, combining ideas from the linear-time algorithms for strongly connected components [55] and dominator trees [4]. Finally, by computing the A-C tree as a preprocessing step, we show how we can improve the time complexity of any SSSP algorithm by running it on the A-C tree. In particular, we improve two state-of-the-art algorithms: Dijkstra’s algorithm [22] with $O(m+n\log(\operatorname{nw}(G)))$ steps, and the recent sparse graph algorithm [19] with $O(m\alpha(n)+m\log^{2/3}(\operatorname{nw}(G)))$ steps, where $\operatorname{nw}(G)\leq n$ is the nesting width of $G$, and $\alpha(n)$ is the extremely slow-growing inverse Ackermann function [54]. Many classes of graphs, such as DAGs, have bounded nesting width, and so on these graphs we obtain SSSP algorithms that run in linear time, a significant step towards beyond-worst-case optimal SSSP algorithms.
@@ 1.4 Related work
Constructing fast shortest path algorithms has been an active area of research since the ‘50s, with pioneering work given by [41, 50, 7, 21, 18], culminating in the work of Fredman and Tarjan [22]. Later work focused on strengthening the assumptions on the inputs. One such approach is to constrain the weights, for example by assuming them to be integers [17, 57, 9, 12] or uniformly bounded [17, 2, 33, 34]. We differ from this line of work since our approach is entirely structural—we assume no restriction on the weights except being non-negative. Other work using structural assumptions include search algorithms for planar graphs [38, 31, 20, 35, 42], minor-free graphs [58], directed acyclic graphs (DAGs) [55], and most recently universal optimality [29]. Generalising the DAG case, [53] showed that SSSP can be solved in time $O(m+n\log k)$ where $k$ is the size of the largest connected component. We obtain this as a special case of our method, because the size of the largest connected component $k$ is always at least $\operatorname{nw}(G)-1$, and is often strictly larger. Finally, there are several recent algorithms using continuous optimisation tools [49, 43, 15, 8, 43, 59]. We work instead directly in a discrete setup to be closer to the original (discrete) problem.
@@
This paper introduces the A-C tree as a hierarchical decomposition of a graph. Similar concepts include the low diameter decomposition (see [9, 12] and references within), which groups together strongly connected sets with respect to the weights (essentially, nodes that are close together should be grouped together), with probabilistic guarantees. In particular, [9, 12] also hierarchically extract acyclic parts using topological order between strongly connected sets to compute shortest path trees faster, but only for integral weights (possibly negative integers). Another line of research is the ‘hierarchy-based’ approach developed by [45, 56, 44, 30]. This approach divides the graph into subgraphs splitting up the original SSSP into smaller SSSPs to search faster. The basic idea is to find subgraphs that are safe to search in greedily (ignoring the remaining graph) up to a certain distance. This idea is similar to how the A-C tree divides the SSSP into smaller problems. However, these hierarchy-type approaches ultimately rely on the weights, whereas our decomposition is entirely structural. Further, it is shown in [45] that any algorithm using this hierarchy method must run in $\Omega(m+n\log n)$ time in the worst case on a certain family of directed graphs. It is straightforward to show that this family has bounded nesting width, and so our algorithm runs in linear time on these graphs.
@@
The A-C tree proposed in this work uses the notion of a dominator tree, introduced by [46] and with the first algorithm given by [39]. Given a graph with a fixed source node $s$, a node $m$ is said to dominate $v$ if every path (from $s$) to $v$ goes through $m$ (Definition 2.2), and the dominator tree is the tree representing this partial order over all nodes (with root $s$). Efficient algorithms for computing the dominator tree include the classical near-linear time algorithm given by [37], and more sophisticated linear time algorithms [4, 13, 14, 24]. In our paper, the dominator tree essentially defines the ‘acyclic’ part of the A-C tree. In this way, the A-C tree can be seen as an extension of the dominator tree, where we also recursively extract strongly connected components. Dominator trees have been used in several applications, such as the design of efficient compilers [1, 16], modelling ecosystems [3], diagnostic test generation [5], 2-vertex connectivity tests and approximations [25, 26], influencing diffusion networks [27], finding strong bridges in graphs [32], diagnosing memory leaks [40], and computing reachability in graphs subject to constraints [47]. However, dominator trees have rarely been applied to shortest path problems, as we do in this paper. One exception is the recent work of [29] which we discussed in the introduction. There, they used the dominator tree in a subroutine to simplify the graph as part of the construction of a universally optimal algorithm for distance ordering in terms of the number of comparison queries. Our usage is similar in flavour, though the dominator tree plays a much more central role in our definitions and results.
@@ 1.5 Outline
The remainder of the paper is as follows. Section 2 presents the preliminaries. Section 3 describes the A-C tree. Section 4 proves that the A-C tree is a minimal-width decomposition. Section 5 presents an algorithm for computing the A-C tree and shows that it has linear-time complexity. This algorithm is then used in Section 6 to solve SSSP through a recursive application of a given SSSP algorithm on the A-C tree, starting with Dijkstra’s algorithm in Section 6.1, followed by an arbitrary SSSP algorithm in Section 6.2. An example illustrating the theory is given in Section 7. Finally, Section 8 concludes the paper. The appendix presents additional results and proofs.